Subject: APO 3.5m users committee minutes, 5/9/11
From: Suzanne Hawley
Submitted: Sat, 14 May 2011 12:51:59 -0700 (PDT)
Message number: 1226
(previous: 1225,
next: 1227
up: Index)
APO 3.5-m Users Committee Phonecon, 5/09/11
Attending: Suzanne Hawley, Mark Klaene, Michael Strauss, Bill Ketzeback,
Scott Anderson, Sean Moran, Jon Holtzman, Cynthia Froning (for John
Bally), Russet McMillan, and Bruce Gillespie
Absent: Al Harper, Remy Indebetouw
**********************************
User feedback and comments from institutional representatives:
o Washington (Scott Anderson) - Scott said that one of the UW users is
interested in an exposure time calculator for the echelle. This was
discussed later in the agenda.
o Colorado (Cynthia Froning) - Cynthia mentioned a user question about a
night log reporting a successful observing night when the primary
instrument scheduled was unavailable. This was discussed later in the
agenda.
o Johns Hopkins (Sean Moran) - Sean had nothing to report.
o New Mexico State (Jon Holtzman) - Jon had nothing to report.
o Princeton (Michael Strauss) - Michael had nothing to report.
o Virginia (Remy Indebetouw) - no report.
o Chicago (Al Harper) - no report.
**********************************
Discussion of telescope/instruments report - Klaene:
Mark mentioned the highlights from the past few weeks. This period was
charactized as windy, dusty, no rain, but with generally good observing
conditions. "Full" closure of the Lincoln Nat'l Forest is expected to be
enacted later this week; access to and operations at APO will be
relatively unaffected, unless an actual fire situation near the
observatory or roads leading to it develops. The telescope has behaved
well, but there is a problem with the Shack-Hartmann instrument for which
there is a remediation plan to replace its camera -- this is an
engineering instrument, so the problem is transparent to the science users
of the telescope. The DIS blue vacuum ion pump failed and was replaced
with a spare. After the ion pump replacement, the cryotiger recovery has
so far not been great, which is typical and is being watched. The Agile
rotator problem was traced to its brake, and was fixed by adjustment. It
was also discovered that the clamps that hold TripleSpec on the rotator
don't work perfectly when the instrument is upside down; a clamp
adjustment and redesign is underway. The echelle camera dewar was pumped.
The NIC-FPS electronics problem is still with us, and we hope this will be
fixed in late June when the instrument goes back to CU. The GFP upgrade
project resumes in early June.
Mark's detailed report for the last month follows:
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights
4/1/11 through 5/5/11
0) Overview
Dry and windy spring weather continues. The USFS enacted stage 2
restrictions, prohibiting all burning and campfires. We expect full
forest closure later this week. This should only affect folks wanting to
walk in the forest. We had a number of severe wind events with speeds
running near 60+ mph. The good news is the intensity and frequency of
these usually drops in May. A few partial nights were lost due to wind
and/or dust. No major fires have occurred near the observatory. A couple
of fires more than 40 miles away did send smoke over us, most of which
dissipated by the end of twilight. Fire prevention work around the
observatory continues as we expect hot and dry weather to continue and the
possibility dry lightning in June.
1) Telescope
The telescope continues to operate routinely with minimal problems during
this period. Some initial summer preventive maintenance work has been
performed. Fritz is working on a new method to map the fiducials on the
az/alt axis to improve blind pointing.
The Shack-Hartmann camera controller failed and a new camera and
controller has been purchased. We will be replacing the Apogee camera
with a new SBIG camera.
2) Instruments
DIS performed well this period until an unplanned Blue camera warm up.
The warm up was due to a failed ION pump which was quickly replaced.
Unfortunately restarting the cryotiger delayed getting the camera back on
line. The instrument has been available with the Red camera only since
the Blue camera went down. There was some confusion the first night on
getting the Red camera only to read out but that was corrected for
subsequent nights. The Blue camera now appears to be working again.
During engineering we removed the slit-viewer assembly to investigate new
mounting options and replacing the camera and optics.
SPICam operated with no new issues.
Agile spontaneous aborts appear to be behind us. The mechanical
installation of a brake on the rotator is substantially complete. A
recent rotator failure has occurred despite our belief that this issue was
fixed. Investigation as to the cause is continuing.
A Triplespec problem has surfaced where the instrument comes off the
kinematic mounts slightly when rotated upside down. No effect on science
has been reported and the instrument safety is not at risk. A temporary
solution to adjust the clamps is in work, and new clamps are being
investigated.
The Echelle's vacuum was serviced during engineering. Chip temperature
has returned to normal. The interorder light is good but slightly degraded
from before pump down. A problem with ecam running during guiding has
been worked on but still is showing up occasionally and work continues on
this.
The NIC-FPS problem of spontaneous reset of the Leach controller when
rotated upside down still exists. Users are advised to start their
observations at specific recommended rotation angles, depending on sky
position, to avoid tracking upside down. See last month's minutes and/or
the instrument manual on the APO website for more information. There is
also a software alert in place that should trigger if an image is bad.
We plan to send the instrument back to CU in late June so the problem can
be addressed during the summer shutdown.
GFP upgrade work will resume in June.
**********************************
Night logs - Hawley:
Recently, the DIS blue camera vacuum ion pumped failed without warning,
and the instrument was taken out of service with only a few hours notice.
The scheduled users were able to invoke a backup program that used SPIcam,
and observations were made. The night log reported that the night was
used, even though the primary program using DIS was not successful. A
user later asked how we reconcile situations like this, i.e., how can the
night be successful if the primary program failed. Suzanne said that the
telescope use and logging tools definitions for a usable night have always
been such that if the telescope collected data, the night was considered
useful (we do not use the word successful in the night logs), and no value
judgement is applied to the science. (Note that this also applies to the
much more frequent case of less than perfect weather - if the weather was
good enough to be open and observing, we log the time as used.) At one
time in the past, we tried a policy where the PIs of the observing
programs were requested to write a short summary assessment of the night's
observing, which would be appended to the night logs. This was
discontinued because the PIs were largely reluctant to provide any
feedback. Without this input from the PIs, we have no tools for
evaluating the quality of the results of observing time when the telescope
is used on the sky.
We want to take this opportunity to remind users that they should always
come to the telescope with a backup program using a different instrument
for the rare case that their primary instrument is unavailable (and in
general with backup programs for poor weather).
In cases of instrument failure, it has been our policy to try to
recompense the affected programs with upcoming OPEN/DD time, which is
usually possible and was done in this case. There followed a discussion
by several committee members that indicated that people feel that our
system seems to work well and is quite flexible.
**********************************
Exposure time calculators - Hawley:
A user has proposed that an exposure calculator be provided for the
echelle. We should consider doing this for all instruments. There is an
exposure graph/table for the echelle on our website, but it is not easily
found and there are some questions about it. But the committee felt that
exposure calculators would be useful tools for observers, and their
development would be a good project for interested members of the
observing community with assistance from APO staff to provide the
necessary instrument parameters. This led to an action item for the
committee:
===> ACTION: User committee members to try to find users who
would be interested in helping develop exposure calculators.
**********************************
Q3 schedule - Klaene, Hawley:
The summer shutdown will run from 5 - 27 July, with the main work being
the realuminization of the primary mirror at Kitt Peak. There will be a
brief engineering period in August for cleanup items. NIC-FIPS will
hopefully be back at APO in late July and available after the shutdown.
There is a UW class trip to APO 11-15 August. Housing is tight at APO
this summer -- a number of people are at the site for the SDSS-III APOGEE
commissioning and other 2.5-m summer work, and NSO housing at Sunspot has
dried up because of the many people being brought there to work on the
ATST project. If any institutions are planning trips to the site this
summer, please contact us as soon as possible to facilitate accomodation
schedules.
**********************************
Final word on logging scripts - Ketzeback:
Bill said that there are now scripts on Newton that match guider images
with the nearest science images, for all instruments. See the link
below, and this will also be announced to apo35-general. We invite
feedback from users.
http://users.apo.nmsu.edu/~bketzeba/ObsTools/Guiders/match.html
**********************************
ReSTAR proposal update - Holtzman:
We received a letter from NOAO that our ReSTAR collaborative proposal
was accepted. We are invited to be part of an NOAO proposal to NSF
for ReSTAR funding that would allow NOAO to lease 3.5m telescope time
for use by the community. The details and extent of the arrangement are
still under discussion.
**********************************
AS3 call for proposals - Gillespie, Hawley:
Although the AS3 call for proposals is mainly targeted to future programs
for the 2.5-m Sloan telescope, some of these proposals may call for
coordinated observing support using the 3.5-m telescope. In these cases,
proposers must contact Suzanne to discuss the feasibility and institutional
arrangements for acquiring access to the 3.5-m prior to submitting their
AS3 proposals.
**********************************
ACTION ITEMS:
[open from previous meeting]:
===> ACTION: User committee members to poll their users to determine who
would be interested in using ARCSAT, willing to help with an upgrade
proposal and its implementation, and for what scientific objectives?
STATUS: Closed. Princeton and JHU showed no interest, while UW,
NMSU, and UC have all indicated some interest. Suzanne will take these
inputs and proceed.
===> ACTION: Bill Ketzeback to look into and suggest fix to permissions
and/or IP issues in TUI related to slit-viewer and guide camera images.
STATUS: Closed. Bill Ketzeback said this has been fixed.
===> ACTION: Users committee members to advise NIC-FPS users about
instrument rotation angle issue and need for vigilance.
STATUS: Closed.
===> ACTION: Remy Indebetouw to get the logging script specifications for
TripleSpec.
STATUS: Closed. Logging scripts have been implemented.
[new actions from this meeting]:
===> ACTION: User committee members to try to find users who would be
interested in helping develop exposure calculators.
**********************************
Next meeting: The next Users Committee phonecon will be on Monday 6 June
at 8:30 AM Pacific Time. The agenda and other materials will be sent to
the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1226 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO